OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER Ted Andrews bought 15 shares of Clark Company stock at $14.875. He sold them a month later at $20.375. What was his profit? Answers: 1.$6.50 2.$49.50 3.$80.50 4.$82.50

OpenStudy (anonymous):

First can you tell me what is profit?

OpenStudy (anonymous):

It's a financial gain

OpenStudy (anonymous):

The answer is 4. $82.50

OpenStudy (anonymous):

It is $1 20.375-14.875=6.50

OpenStudy (anonymous):

opps i meant #1

OpenStudy (anonymous):

Just do 14.875 x 15. Which is 223 dollars. Rounded. Then multiply 20.375 by 15. Which is 305 dollars. 305-223 = 82 Dollarss. So #4 is the right awnser :D

OpenStudy (anonymous):

#TakingMedal :D

OpenStudy (anonymous):

noo you are taking the amount he bought the share and the amount he sold it at and subtrating it

OpenStudy (anonymous):

Exactly. He bought 15 shares of it. So 15 gets multiplied by how much he sold it for. :D

OpenStudy (anonymous):

because he sold them at a higher price so you would get 6.50

OpenStudy (anonymous):

IT's 6.50, and you multiply it by 15 because he bought 15 of them

OpenStudy (anonymous):

SO the answer would be 82.50$

OpenStudy (anonymous):

Exactly. So I deserve the medal :D

OpenStudy (anonymous):

no he saying he sold all of them for that price it does not say he sold each of them for twenty dollar

OpenStudy (anonymous):

He sold each of them for 20 dollars. He bought each of them for 14 dollars. The answer is #4. Just trust me.

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

9 th it doesnt say that

OpenStudy (anonymous):

14.875 times 15 = 223.125 (How much he payed for them all together) 20.375 times 15 = 305.625 (how much he sold them for) 305.625- 223.125= $82.5

OpenStudy (anonymous):

So my answer is correct. Sammy224 it could be confusing but you have to think about it. You can't just answer the question without thinking. :D Just saying.